# Generated Data with Fake Privacy: Hidden Dangers of Fine-tuning Large Language Models on Generated Data

### v1

@@ 4 Privacy Risks on Self-Instruct Tuning
To reduce the cost of instruction-tuning, Wang et al. [34] proposed the ’Self-Instruct’ method, which has since been widely adopted in training various large language models (LLMs), such as Alpaca [31]. In this section, we apply self-instruct tuning to legal LLMs, a popular instruction-tuning task where the training data often contains sensitive information. After completing the tuning process, we investigate the potential privacy risks associated with the resulting legal chatbot using the MIA attack. Finally, we explore the relationship between privacy and utility in self-instruct models following the pipeline drawn in Figure 6.
@@ 4.1 Experiment Settings
In this section, we continue to utilize the Pythia models as targets for fine-tuning and evaluation, consistent with the approach outlined in Section 3. However, we replaced the Pythia-12b model with Llama-3-8B-Instruct as the generator, due to the latter’s superior ability to follow the given context and produce more coherent and relevant data. Following the pipeline shown in Figure 6, detailed information on data generation (including random sampling, instruction generation, and input-out generations), model fine-tuning, and evaluation procedures is provided below.
@@ 4.1.1 Data Generation
Compared to the data used for supervised fine-tuning in Section 3, the data structure for instruction-tuning is more complex, as it typically includes task descriptions, task-related inputs, and the corresponding outputs. To generate such data samples, self-instruct tuning involves querying a generator using predefined contexts, denoted as seed tasks. These seed tasks contain task descriptions along with associated input-output pairs. Guided by these seed tasks, the generator can produce the necessary data samples for instruction-tuning.
To generate fine-tuning data for legal language models, we construct seed tasks with corresponding input-output pairs, following the pipeline outlined by Wang et al. [34]. We create $75$ input-output pairs in total for 64 seed tasks. The seed tasks are constructed with manual work and their inputs are selected from the test split of FreeLaw, a dataset included in Pythia’s pre-training set of legal documents. The details are listed as follows.
All $64$ task descriptions of the legal LLMs are listed in the Appendix. Then we use the instructions from the seed dataset for 3-shot prompting on the Llama3-Instruct-8B and collected 4000 new instructions. For bootstrapping, we have used 4-shot prompting. Finally, we filter the examples with low quality, whose input and outputs are not explicitly presented, and then get the generation data for self-instruct tuning.
4.1.2 Other Details
After obtaining the generated data related to the legal tasks, we do QLoRA [13] fine-tuning for Pythia-6.9b model with Adam optimizer [20] for $1$ epoch with $64$ rank, $1/4$ scale factor, $0.05$ LoRA dropout rate, and batch size of $8$. Like Section 3, we also search the training hyper-parameters to evaluate the worst cases of privacy leakage. Since data size and data quality will greatly influence the performance of the obtained legal LLMs, we also search the data size and temperature for self-instruct tuning. Details of the search space are listed in Table 2
Table 2: Hyper-parameters used for Pythia’s self-instruct tuning.
@@ 4.1.2 Other Details
After merging the adapters with the appropriate base model and converting the parameters into 16-bit, we can get the fine-tuned legal LLM based on Pythia.
The finetuning process is done for four learning rates, three dataset sizes, and eight temperatures for the generator model, which are presented in the table below.
@@ 4.1.3 Evaluation Metrics
  • CaseHOLD dataset provides 53k cases, each of which includes five multiple-choice options for the relevant legal holding.
  • SSLA is a popular subset of the widely used LegalBench [17], including 1038 samples on "plaintiff", 1016 samples on "individual defendants", and 1234 samples on"company defendants".
After evaluating the model’s utility, we then construct validation sets for privacy. We randomly choose $100$ samples from FreeLaw’s training set as members and $100$ samples from FreeLaw’s test set as non-members for the membership inference attack. For finetuned models, both members and non-members are placed in alpaca prompts while for base models, they are used in raw form.
Utility Evaluation For the utility’s measurement, we first query the fine-tuned and base models with the prompts in the CaseHOLD datasets and get the responses. The responses with explicitly “holding number” are considered valid. Then we count all correct answers in the valid responses with the ground truth target and calculate the accuracy of each model. After that, we feed LLMs with queries in SSLA and then calculate the similarity score [27] of the generated responses and the ground truth answers. If the fuzz score is larger than $80$ we count it as accurately answering the legal questions in SSLA.
@@
The attack performances of base models are used as baselines for fine-tuned models. For the MIA on the instruction-tuned models, we place member and non-member documents inside the alpaca prompt template [31] as illustrated in Figure 7. For base models, on the other hand, we use their original format. This is because the fine-tuned models have been trained on alpaca format whereas base models have been trained on raw texts. An important remark is that the attacker is free to choose the best template for maximizing the attack performance.
Figure 7: Member and Non-Member sample construction process for Membership Inference Attack on instruction-tuned models.
@@ 4.2.1 Utility Evaluation
After self-instruct tuning, we first evaluate the model’s utility to demonstrate that our self-instruct tuning is conducted properly. Firstly, we use the CaseHOLD tasks for evaluation, which include 5-choice legal questions on appropriate holdings for the given case. Since the CaseHOLD tasks aim only for multiple-choice, it cannot properly evaluate the model’s performance as a chatbot. Therefore, we also use the SSLA tasks for evaluation. It consists of purely generation tasks and the performance is assessed with respect to the accuracy and conciseness of the answers. The results for self-instruct setting and the base model are listed in Table 3
@@
From the table, one can see that we effectively perform the Self-Instruct method. The average improvement on CaseHOLD tasks is doubled. As for SSLA tasks, Pythia 6.9b also achieves more than $20\%$ improvements after the self-instruct tuning. Furthermore, we also draw an example of SSLA in Figure 8 to demonstrate the utility improvement after the self-instruct tuning. From the figure, one can see that the LLMs can follow the instructions on law cases to find the plaintiffs in given except after the self-instruct tuning while the base model only replies to some related but incorrect sentences.
Figure 8: A case of SSLA’s tasks and the response from the base Pythia 6.9b and its self-instruct tuned version.
@@ 4.2.2 Privacy Evaluation
To measure the model’s privacy risks after the self-instruct tuning, we conduct four MIA methods described in Section 4.1.3. We first listed the best ROC-AUC score of Pythia 6.9b after the self-instruct tuning across various hyper-parameter settings in Table 4.
@@
Table 4: Highest ROC-AUC of base and self-instruct tuned Pythia 6.9b across different hyper-parameter settings under different MIA methods.
From the results, one can see that the base Pythia 6.9b can be considered to be safe under different membership inference attacks, as the ROC-AUC score for different methods is around $0.5$. It demonstrates that all these MIA methods perform like random guesses on the base model, which is consistent with former research [15]. However, we also find that the ROC-AUC score for the Pythia 6.9b model after the self-instruct tuning improves a lot, over $40\%$ for each attack. Such results demonstrate that self-instruct tuning can make Pythia 6.9b greatly vulnerable to MIA and lead to serious privacy risks on the model’s pre-training data.
Figure 9: The distributions of the ROC-AUC score when conducting MIA on models tuned with all the hyper-parameter settings stated in section 4.1.2. The red line shows the ROC-AUC score of the base Pythia-6.9b model. Blue bins represent the number of fine-tuned models. The x-axis denotes the ROC-AUC score.
Apart from the self-instruct tuned model with the top privacy risks, we also report the distributions of the ROC-AUC score for models with all the hyper-parameter settings. The results are shown in Figure 9. From the figure, one can see that the ROC-AUC scores for all the models are higher than the base models with a large margin. The worst improvement is still larger than the $20\%$ increment compared with the base model under different attacks. The results demonstrate that the self-instruct tuning may cause privacy risks in nearly all cases.
4.3 Ablations Studies
In this section, we try to explore the key factors that influence the models’ privacy the most after the self-instruct tuning. We explore the key factors stated in 4.1.2, including the temperature, learning rate, and datasets.
4.3.1 Temperatures for the Data Generation
Different temperature settings for the generator will related to different quality of the input-output examples. Therefore, we plot the averaged ROC-AUC score for different MIA methods on Pythia 6.9b, which is fine-tuned on the self-instruct data generated with different temperatures. As for other hyper-parameters, we choose the learning rate to be $2\times 10^{-4}$ and the data size to be $250$. It is the same setting for the self-instruct tuned model with the highest privacy risks.
@@ 4.3.1 Temperatures for the Data Generation
From Figure 10, one can see that different temperatures do not influence the ROC-AUC score much. The highest ROC-AUC score is 0.71 with a temperature near 00 while the lowest ROC-AUC score is 0.69 with a temperature equal to $1.4$. However, there exists a consistent trend that when the temperature increases, the averaged ROC AUC score continually goes smaller. This is because the high temperature of the generator model will make the generated data less similar to the original pre-training dataset. Therefore, LLM’s memories of the training data and the ROC-AUC score will be lower. Therefore, we recommend using a larger temperature to alleviate the privacy risks.
4.3.2 Learning Rate for the Self-Instruct Tuning
According to the former analysis, the learning rate is a key factor that influences both the model’s memorization and the potential privacy risks. Therefore, we conduct experiments to explore the learning rate’s effect on models after the self-instruct tuning. To investigate the effect of learning rate in isolation, we fix the generator temperature to $0.6$ and dataset size of $250$. The results are drawn in Figure 11.
Figure 11: The averaged ROC-AUC score of different MIA methods conducted on self-instruct tuned models with different learning rates.
From the figure, one can see that the ROC-AUC scores substantially correlate with the learning rate increasing, the improvement is more than $20\%$ when changing the learning rate from $2\times 10^{-6}$ to $2\times 10^{-3}$. A similar phenomenon is also observed when training with real data by many previous studies [11, 9, 19]. The possible reason for such improvement is the larger learning rate makes LLMs better memorize the fine-tuning data and also activates the memorization of the pre-training datasets. Therefore, the MIA methods can perform better in such scenarios.
Apart from learning rates’ influence on privacy, we also compare the model’s utility fine-tuned with different learning rates. The results are listed in Table 5. Combined with the results in Figure 11, one can see that a larger learning rate will enhance both the utility and the privacy risks, as the models fit better in such settings. Moreover, we also find that using a smaller learning round $10^{-4}$ can reduce the AUC ROC’s performance with a good performance.
@@ 4.3.2 Learning Rate for the Self-Instruct Tuning
4.3.3 Data Size for the Self-Instruct Tuning
Beyond the learning rate, the size of the self-instruct dataset may also impact the privacy of LLMs after self-instruct tuning. To investigate this, we conduct experiments to examine how dataset size affects MIA performance. We plot the ROC-AUC scores for different MIA methods across models fine-tuned with datasets ranging from 250 to 4,000 samples, as shown in Figure 12. These experiments are conducted with three different learning rates: $2\times 10^{-5}$, $2\times 10^{-4}$, and $2\times 10^{-3}$.
@@ 4.3.3 Data Size for the Self-Instruct Tuning
The results indicate that all MIA methods exhibit similar trends when varying learning rates and datasets. A smaller learning rate notably enhances the ROC-AUC score as the dataset size increases. This is particularly evident for the reference-based attack, loss attack, and Min-k attack, where the ROC-AUC score improves by $10\%-20\%$ when the dataset size is scaled from 250 to 4,000. However, with larger learning rates, the differences between models fine-tuned with varying dataset sizes are less pronounced. This may be because models trained with smaller learning rates require more data to converge, while larger learning rates enable models to quickly memorize patterns similar to the original training samples in the self-instructed data, resulting in higher ROC-AUC scores after fine-tuning. Nevertheless, due to differences between the self-instructed data and the original pre-training data, the ROC-AUC score only increases to around 0.7-0.8. Overall, the findings suggest that both larger learning rates and dataset sizes can heighten privacy risks up to a certain threshold. Therefore, we recommend using a slightly smaller learning rate and dataset size to manage these risks effectively.
4.4 Takeway Messages
In this section, we assess the privacy risks of LLMs on their pre-training datasets following self-instruct tuning. Using the example of a legal chatbot, we adopt the self-instruct pipeline to train a legal LLM based on Pythia 6.9b and then evaluate both the model’s utility and privacy. The results indicate that self-instruct tuning can significantly increase privacy risks associated with the pre-training dataset, FreeLaw, with over a $40\%$ improvement in ROC-AUC scores across various MIA methods. Additionally, our experiments reveal that the learning rate and dataset size are critical factors influencing privacy risks. Higher learning rates and larger datasets make the fine-tuned model more susceptible to membership inference attacks. Consequently, we recommend opting for a slightly smaller learning rate and dataset size during training to better safeguard privacy.

### v2

@@ 4 Privacy Risks on Self-Instruct Tuning
To reduce the cost of instruction-tuning, Wang et al. [52] proposed the ’self-instruct’ method, which has since been widely adopted in training various LLMs, such as Alpaca [48]. In this section, we apply self-instruct tuning to legal LLMs, a popular instruction-tuning task where the training data often contains sensitive information. After completing the tuning process, we investigate the potential privacy risks associated with the resulting legal chatbot using the MIA attack. Finally, we explore the relationship between privacy and utility in self-instruct models following the pipeline drawn in Figure 8.
@@ 4.1 Experiment Settings
In this section, we use Pythia models as fine-tuning targets and for evaluation, aligning with the first experimental methodology in Section 3. However, we replaced the Pythia-12b model with Llama-3-8B-Instruct as the generator, due to the latter’s superior ability to follow the given context and produce more coherent and relevant data. Following the pipeline illustrated in Figure 8, we provide detailed information on the data generation process (including random sampling, instruction creation, and input-output generation), as well as the procedures for model fine-tuning and evaluation.
@@ 4.1.1 Data Generation
Compared to the data used for supervised fine-tuning in Section 3, the data structure for instruction-tuning is more complex, as it typically includes task descriptions, task-related inputs, and the corresponding outputs. To generate such data samples, self-instruct tuning involves querying a generator using predefined contexts, denoted as seed tasks. These seed tasks contain task descriptions along with associated input-output pairs. Guided by these seed tasks, the generator can produce the necessary data samples for instruction tuning.
To generate fine-tuning data for legal language models, we construct seed tasks with corresponding input-output pairs, following the pipeline outlined in [52]. We create $75$ input-output pairs in total for 64 seed tasks. The seed tasks are manually crafted with inputs selected from FreeLaw’s test split, part of Pythia’s pre-training legal dataset. Then, we use the instructions from the seed dataset for 3-shot prompting on the Llama3-Instruct-8B and collect 4000 new instructions. For bootstrapping, we use 4-shot prompting. Finally, we filter out low-quality examples where inputs and outputs are not explicitly defined, resulting in a refined dataset for self-instruct fine-tuning.
4.1.2 Fine-tuning Details
After obtaining the generated data related to the legal tasks, we perform QLoRA [19] fine-tuning for Pythia-6.9b model with Adam optimizer [30] for $1$ epoch with $64$ rank, $1/4$ scale factor, $0.05$ LoRA dropout rate, and batch size of $8$. Similar to Section 3, we also explore various training hyperparameters to assess the worst-case scenarios of privacy leakage. Since data size and data quality will greatly influence the performance of the obtained legal LLMs, we also search the data size and temperature for self-instruct tuning. Details of the search space are listed in Table 2.
@@ 4.1.2 Fine-tuning Details
Table 2: Hyperparameters for Pythia’s self-instruct tuning.
After merging the adapters with the appropriate base model and converting the parameters into 16-bit, we get the fine-tuned legal LLM based on Pythia. The finetuning process is done for four learning rates, three dataset sizes, and eight temperatures for the generator model as summarized in Table 2.
@@ 4.1.3 Evaluation Metrics
CaseHOLD dataset comprises 53k legal cases, each accompanied by five multiple-choice options corresponding to the relevant legal holding.
SSLA is a subset of the widely used LegalBench [26], including 1038 samples on "plaintiff", 1016 samples on "individual defendants", and 1234 samples on "company defendants".
After evaluating the model’s utility, we construct validation sets for privacy. We randomly choose $100$ samples from FreeLaw’s training set as members and $100$ samples from FreeLaw’s test set as non-members for the membership inference attack. For finetuned models, both members and non-members are placed in Alpaca prompts while for base models, they are used in raw form.
Utility Evaluation For utility measurement, we first query the fine-tuned and base models with the prompts in the CaseHOLD datasets and get the responses. The responses with explicitly “holding number” are considered valid. Then, we count all correct answers in the valid responses with the ground truth target and calculate the accuracy of each model. After that, we feed LLMs with queries in SSLA and then calculate the similarity score [44] of the generated responses and the ground truth answers. If the fuzz score is larger than $80$ we count it as accurately answering the legal questions in SSLA.
@@
The attack performances of base models are used as baselines for fine-tuned models. For the MIA on the instruction-tuned models, we place member and non-member documents inside the Alpaca prompt template [48] as illustrated in Figure LABEL:fig:mia-construction-diagram, Appendix A. For base models, we use their original format. This is because the fine-tuned models have been trained on Alpaca format whereas base models have been trained on raw texts. An important remark is that the attacker is free to choose the best template for maximizing the attack performance.
Member/Non-Member Partitioning The member samples for MIA are randomly chosen from the train split of FreeLaw as in the work of Duan et al. [21]. In parallel, the non-member samples are randomly chosen from the test split of FreeLaw. The samples in two groups share the same style and content characteristics. This poses a real challenge for MIA in distinguishing members based only on their intrinsic properties, thus providing a robust evaluation of privacy risks.
@@ 4.2.1 Utility Evaluation
We evaluate the model’s utility to demonstrate that our self-instruct tuning is properly conducted. Since the CaseHOLD tasks aim only for multiple-choice, it cannot properly evaluate the model’s performance as a chatbot. Therefore, we also use the SSLA tasks for evaluation. It consists of purely generation tasks and the performance is assessed with respect to the accuracy and conciseness of the answers. The results for the self-instruct setting and the base model are listed in Table 3.
@@
We observe that we effectively perform the Self-Instruct method. The average improvement on CaseHOLD tasks is doubled. As for SSLA tasks, Pythia 6.9b also achieves more than $20\%$ improvements after the self-instruct tuning. Furthermore, we draw an example of SSLA in Figure 9 to demonstrate the utility improvement after self-instruct tuning.
Figure 9: A case of SSLA’s tasks and the response from the base Pythia 6.9b and its self-instruct tuned version. The base model returns an irrelevant and generalized reply, while the fine-tuned model returns a direct and privacy-sensitive reply that satisfies the requirement of the task.
@@ 4.2.2 Privacy Evaluation
To measure the model’s privacy risks after the self-instruct tuning, we conduct four MIA methods described in Section 4.1.3. We first listed the best ROC-AUC score of Pythia 6.9b after the self-instruct tuning across various hyperparameter settings in Table 4.
@@
Table 4: Highest ROC-AUC of base and self-instruct tuned Pythia 6.9b across different hyperparameter settings under different MIA methods.
From the results, we observe that the base Pythia 6.9b can be considered to be safe under different membership inference attacks, as the ROC-AUC score for different methods is around $0.5$. It demonstrates that all the MIA methods perform similarly to random guessing on the base model, aligning with findings from previous research [21]. However, we also find that the ROC-AUC score for the Pythia 6.9b model after the self-instruct tuning improves over $40\%$ for each attack. Such results demonstrate that self-instruct tuning can make Pythia 6.9b greatly vulnerable to MIA and lead to serious privacy risks on the model’s pre-training data.
Figure 10: The distributions of the ROC-AUC score for MIA on models tuned with all the hyperparameter settings stated in Section 4.1.2. The red line shows the ROC-AUC score of the base Pythia-6.9b model. Blue bars represent the number of fine-tuned models. The x-axis denotes the ROC-AUC score.
Apart from the self-instruct tuned model with the top privacy risks, we also report the distributions of the ROC-AUC score for models with all the hyperparameter settings. The results are shown in Figure 10. We demonstrate that the ROC-AUC scores for all the models are higher than the base models with a large margin. The worst improvement is still larger than the $20\%$ increment compared with the base model under different attacks. The results reflect that the self-instruct tuning may cause privacy risks in nearly all cases.
4.3 Ablation Studies
In this section, we explore the key factors that influence the models’ privacy leakage after the self-instruct tuning. We explore the key factors stated in Section 4.1.2, including the temperature, learning rate, and datasets.
The impact of temperature. Different temperature settings for the generator affect the quality of the input-output examples, leading to variations in their relevance and diversity. Therefore, we plot the averaged ROC-AUC score for different MIA methods on Pythia 6.9b, which is fine-tuned on the self-instruct data generated with different temperatures. As for other hyperparameters, we choose the learning rate to be $2\times 10^{-4}$ and the data size to be $250$. It is the same setting for the self-instruct tuned model with the highest privacy risk.
@@ 4.3 Ablation Studies
Figure 11 illustrates that temperature variations have minimal effect on the ROC-AUC score, with the highest score of 0.71 observed near a temperature of 0 and the lowest score of 0.69 at a temperature of 1.4. However, a consistent trend emerges where an increase in temperature is associated with a gradual decline in the averaged ROC-AUC score. This occurs because a higher temperature in the generator model reduces the similarity between the generated data and the original pre-training dataset. Thus, LLM’s memories of the training data and the ROC-AUC score will be weaker and we recommend using a larger temperature to alleviate the privacy risks.
The impact of learning rate. To investigate learning rate’s effect in isolation, we fix the generator temperature to $0.6$ and dataset size to $250$. The results are drawn in Figure 12.
Figure 12: The averaged ROC-AUC score of different MIA methods conducted on self-instruct tuned models with various learning rates.
We observe a substantial correlation between ROC-AUC scores and increasing learning rates, with an improvement exceeding $20\%$ when the learning rate is adjusted from $2\times 10^{-6}$ to $2\times 10^{-3}$. A similar phenomenon is also observed when training with real data by previous studies [16, 15, 28]. A possible reason for such improvement is the larger learning rate makes LLMs better memorize the fine-tuning data and also activates the memorization of the pre-training datasets. Therefore, the MIA methods can perform better in such scenarios.
Apart from learning rates’ influence on privacy, we also compare the model’s utility fine-tuned with different learning rates. The results are listed in Table 5. Combined with the results in Figure 12, we see that a larger learning rate will enhance both the utility and the privacy risks, as the models fit better in such settings. Moreover, we also find that using a smaller learning round $10^{-4}$ can reduce the AUC ROC’s performance with a good performance.
@@
The impact of dataset size. We explore how self-instruct dataset size impacts LLM privacy through experiments analyzing its effect on MIA performance. We plot the ROC-AUC scores for different MIA methods across models fine-tuned with datasets ranging from $250$ to $4,000$ samples, as shown in Figure 13. These experiments are conducted with three different learning rates: $2\times 10^{-5}$, $2\times 10^{-4}$, and $2\times 10^{-3}$.
@@
The results indicate that all MIA methods display similar trends across different learning rates and datasets. A smaller learning rate notably enhances the ROC-AUC score as the dataset size increases. This is particularly evident for the reference-based attack, loss attack, and Min-k attack, where the ROC-AUC score improves by $10\%-20\%$ when the dataset size is scaled from $250$ to $4,000$. However, with larger learning rates, the differences between models fine-tuned with varying dataset sizes are less pronounced. This may be because models trained with smaller learning rates require more data to converge, while larger learning rates enable models to quickly memorize patterns similar to the original training samples in the self-instructed data, resulting in higher ROC-AUC scores after fine-tuning. Nevertheless, due to differences between the self-instructed data and the original pre-training data, the ROC-AUC score only increases to around 0.7-0.8. Overall, the findings suggest that both larger learning rates and increased dataset sizes can amplify privacy risks up to a certain threshold. Therefore, we recommend using a slightly smaller learning rate and dataset size to manage these risks effectively.
4.4 Discussion
In this section, we assess the privacy risks of LLMs on their pre-training datasets following self-instruct tuning. Using the example of a legal chatbot, we adopt the self-instruct pipeline to train a legal LLM based on Pythia 6.9b and evaluate both the model’s utility and privacy. We observe that self-instruct tuning can substantially increase privacy risks for the pre-training dataset, FreeLaw, with over a $40\%$ improvement in ROC-AUC scores across various MIA methods. Additionally, our experiments reveal that the learning rate and dataset size are critical factors influencing privacy risks. Higher learning rates and larger datasets make the fine-tuned model more susceptible to membership inference attacks. Consequently, we recommend opting for a slightly lower learning rate and dataset size during training to safeguard privacy. While the suggested mitigations, such as using a lower learning rate (Section 3) or higher temperatures (Section 4), help to reduce the privacy risks, they do not entirely eliminate them, underscoring the need for further research and complementary approaches to fully address these vulnerabilities.
In addition, former methods such as differential privacy (DP) [36, 10], data anonymization [25], and data augmentation [20] can be used to alleviate the consequences of privacy leakage. DP can reduce memorization and mitigate membership inference, PII extraction, and model inversion risks by introducing noise [10]. Moreover, DP enables risk estimates through theoretical guarantees [7]. Data augmentation trains the model on multiple closely related data points rather than a single instance, potentially mitigating the impact of memorization [54]. Yet, it is important to note that certain MIA implementations can leverage the structural similarities among augmented data points to enhance the success rate of attacks [35]. Although data anonymization may not directly reduce memorization, it can render potential leakages less harmful and provide better safety standards, especially for PII attacks [25]. Future research should explore hybrid approaches that combine these techniques to enhance data privacy without significantly compromising model utility.